Two angles are supplementary. One is 56 degrees more than 3 times the other. Find the measures of the angles
Answer by stanbon(75887)

Two angles are supplementary. One is 56 degrees more than 3 times the other. Find the measures of the angles
------
Angle: x
Supplementary Angle: 180-x
-------------------------------
Equation:
x = 3(180-x)+56
-------
x = 3*180-3x+56
---
4x = 596
---
x = 149
180-x = 31
==============
Cheers,
Stan H.